OpenStudy (anonymous):

Solve: x + 9 = 5(4x – 2) Select one: a.  b. –1 c. 1 d. 

OpenStudy (anonymous):

What is the answer

OpenStudy (anonymous):

I will really appreciate it if someone gives it to me

OpenStudy (anonymous):

A is 11/9 B is -1/19

OpenStudy (priyar):

x + 9 = 20x - 10 [multiplying 5 inside] ok?

OpenStudy (priyar):

@Jesus98 ?

OpenStudy (priyar):

x + 9 = 5(4x – 2) x + 9 = 20x - 10 x-x+9 = 20x-x-10 subtacting x on both sides 9+10 = 19x adding 10 on both sides 19= 19x

OpenStudy (priyar):

so x=?

OpenStudy (priyar):

@Jesus98 pls respond..any doubts so far?

OpenStudy (priyar):

19x=19 now divide 19 on both sides..thats it.

OpenStudy (priyar):

so x =?

OpenStudy (whpalmer4):

\[x + 9 = 5(4x – 2)\]Rules of solving equations are that you can do pretty much anything, as long as you do it to both sides. \[x+9 = 5(4x-2)\]first we use the distributive property to get rid of the parentheses:\[x+9=5*4x+5(-2)\]\[x+9=20x-10\] now we want to get \(x\) by itself on the left side, and everything else on the right. We start by subtracting \(9\) from each side: \[x+9-9 = 20x-10-9\]\[x=20x-19\]now we subtract \(20x\) from each side: \[x-20x=20x-20x-19\]\[x-20x=-19\]\[1x-20x=-19\]\[x(1-20) = -19\]\[x(-19) = -19\]divide both sides by \(-19\)\[\frac{x(-19)}{-19} = \frac{-19}{-19}\]\[x=1\] Important step: we check our answer by plugging it into the original formula and make sure it is true: \[x+9=5(4x-2)\]\[1+9 = 5(4(1)-2)\]\[10=5(4-2)\]\[10=5(2)\]\[10=10\checkmark\]that made a true statement, so our answer is correct.

OpenStudy (anonymous):

Thanks

OpenStudy (anonymous):

I got one more for you

OpenStudy (anonymous):

Whpalmer4

OpenStudy (anonymous):

5p – 14 = 8p + 4

OpenStudy (carra.scott):

C. 1

OpenStudy (anonymous):

c

OpenStudy (whpalmer4):

\[5p-14=8p+4\] subtract \(5p\) from both sides. subtract \(4\) from both sides. divide both sides by the coefficient of \(p\) (the number in front of it — in \(23c\), the coefficient is \(23\)) again, check your answer for correctness by plugging it into the equation and verifying a true statement results.

OpenStudy (anonymous):

x=1

OpenStudy (anonymous):

p=-6

OpenStudy (anonymous):